OpenStudy (anonymous):

3/4x-8=8-3/4x (3/4 are fractions) plz include answer

OpenStudy (anonymous):

bring the x-variables to one side and constants to the other. 3/4x + 3/4x = 8+8 3/2x = 16 x=32/3

OpenStudy (anonymous):

\[X=\frac{ 32 }{ 3 }\]

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

then reduce

OpenStudy (anonymous):

10 2/3 i beleive

OpenStudy (anonymous):

x=10 \[\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

or 32/3

OpenStudy (anonymous):

thx